A 180-pound male has a VO2 of 7.3 METs and recently has had leg surgery. He has been ordered to being arm cycling 3 days/week at a speed of 60 rpm. How much resistance should be set on the ergometer?
Answer #1

Weight of the patient - 180lbs = 81kg

VO2 - 7.3 METs( 1MET = 3.5 ml/kg/min ) = 2069.5 ml/min

Ordered arm cycling 3days /week at a speed of 60rpm

Resistance set on the ergometer :-

Using formula :-

VO2 = (1.8 ×work rate in kgm /min )/ body weight in kg +7

= VO2 /kg min = 2069.5/81 = 25.5 ml/min

RPM = 60× 6 = 360 m/min× 81 = 29160 kgm/min

Putting values into the formula :-

= 25.5 ml/kgmin =( 1.8 × (F) 360/81 ) +7

= 25.5 - 7 = 1.8× 360F / 81

= 18.5× 81 = 1.8× 360F

= 1498.5 = 648F

F = 1498.5/ 648

= 2.31 kg is the resistance